### v1

# DiffUHaul: A Training-Free Method for Object Dragging in Images

Abstract.
Text-to-image diffusion models have proven effective for solving many image editing tasks. However, the seemingly straightforward task of seamlessly relocating objects within a scene remains surprisingly challenging. Existing methods addressing this problem often struggle to function reliably in real-world scenarios due to lacking spatial reasoning. In this work, we propose a training-free method, dubbed DiffUHaul, that harnesses the spatial understanding of a localized text-to-image model, for the object dragging task. Blindly manipulating layout inputs of the localized model tends to cause low editing performance due to the intrinsic entanglement of object representation in the model. To this end, we first apply attention masking in each denoising step to make the generation more disentangled across different objects and adopt the self-attention sharing mechanism to preserve the high-level object appearance. Furthermore, we propose a new diffusion anchoring technique: in the early denoising steps, we interpolate the attention features between source and target images to smoothly fuse new layouts with the original appearance; in the later denoising steps, we pass the localized features from the source images to the interpolated images to retain fine-grained object details. To adapt DiffUHaul to real-image editing, we apply a DDPM self-attention bucketing that can better reconstruct real images with the localized model. Finally, we introduce an automated evaluation pipeline for this task and showcase the efficacy of our method. Our results are reinforced through a user preference study.
Object Draggining, Image Editing
journal: TOGsubmissionid: 0ccs: Computing methodologies Machine learningccs: Computing methodologies Computer graphics
Figure 1. DiffUHaul: Given a real image with multiple objects (e.g., a cat and a rock), our method is able to seamlessly drag each of the objects to an arbitrary location within the image while preserving the foreground and background appearance.
Project page is available at: https://omriavrahami.com/diffuhaul/
1. Introduction
Think about a digital artist who recently employed an advanced generative model to craft an image featuring a Persian cat alongside a rock, as in Figure 1. All that is needed for their creation to achieve perfection is for the cat (or the rock) to be moved slightly. Despite the conceptual simplicity of such a task, seamlessly dragging objects in an image is surprisingly challenging for current generative image editing methods (Brooks et al., 2023; Hertz et al., 2022). In this work, we propose a novel training-free solution for this scenario.
Current methods that tackle this problem rely on time-consuming LoRA training per image (Shi et al., 2023), training a designated model on a large dataset (Chen et al., 2023a; Yang et al., 2022) or utilizing classifier-free guidance (CFG) with specific objectives (Mou et al., 2023, 2024; Epstein et al., 2023). However, these methods are not robust and struggle to operate reliably in a real-world setting. For example, as can be seen in Figure 2, DiffEdit (Mou et al., 2024) suffers from artifacts of traces of the puppy in its original location, while our method demonstrates a more robust behavior.
Recently, several localized text-to-image models were developed by the community that add spatial controllability to the task of text-to-image generation (Li et al., 2023; Avrahami et al., 2023c; Yang et al., 2023; Zheng et al., 2023; Zhang et al., 2023; Nie et al., 2024). A natural question is then whether the localized understanding of the 2D pixel world in such models can be harnessed for the task of object dragging. Hence, we examine the disentanglement properties of such models, and propose a series of modifications that allow them to serve as a backbone for drag-and-drop movement of objects within an image. Specifically, we use the recently introduced BlobGEN (Nie et al., 2024) model, and demonstrate that its spatial understanding can enable significantly more robust object dragging without requiring fine-tuning or training.
DiffEdit
Refer to caption Refer to caption Refer to caption Refer to caption
DiffUHaul (ours)
Refer to caption Refer to caption Refer to caption Refer to caption
Input Dragging the puppy from left to right ($\rightarrow$)
Figure 2. Object Dragging Robustness. When dragging a puppy in a complex environment (particularly with its reflection in the water and ripples nearby) to different locations along from left to right, previous method DiffEdit (Mou et al., 2024) struggles with the editing traces left in its original location, while our method demonstrates a more robust behavior.
In pursuit of our solution, we begin by revealing an entanglement problem in the localized text-to-image models, through which the prompt-based localized controls of different image regions interfere with each other. We trace the root cause to the commonly used Gated Self-Attention layers (Li et al., 2023), where each individual layout embedding are free to attend to all the visual features. We propose an inference-time masking-based solution, named gated self-attention masking, and show that improving the model disentanglement leads to better object dragging performance.
Next, specially for the object dragging task, we first adopt the commonly-used self-attention sharing mechanism (Cao et al., 2023) to preserve the high-level object appearance. To better transfer the fine-grained object details from source images to target images and better harness spatial understanding of the model, we propose a novel soft anchoring mechanism: in early denoising steps, which control the object shape and scene layout in an image, we interpolate the self-attention features of the source image and those of the target image with a coefficient relative to the diffusion time step. This process promotes a smooth fusion between the target layout and source appearance. Then, in later denoising steps, which control the fine-grained visual appearance in an image, we update the interpolated attention features from the corresponding features in the source image via the nearest-neighbor copying.
To adapt our method to real-image editing, we further require an inversion solution that is compatible with the localized method. We find that the standard DDIM inversion (Song et al., 2020) struggles to reconstruct the image faithfully, even when not using classifier-free guidance (Ho, 2022). Hence, we propose a simple DDPM self-attention bucketing technique that adds noise to the reference image independently in each diffusion step, and uses the noisy images to extract the self-attention outputs as the source attention features. This DDPM bucketing does not accumulate reconstruction errors along the denoising process and preserves details for real images.
Finally, we offer automatic metrics for our problem to assess different aspects of the editing operations, and use them for an extensive comparison that demonstrates the effectiveness of our method over the baselines. In addition, we conduct a user study and show that our method is also preferred by human evaluators.
In summary, our contributions are: (1) we show that the spatial understanding of a localized text-to-image model can be effectively harnessed to tackle the object dragging task, (2) we reveal an entanglement problem in the gated self-attention layers and offer an inference-time solution, (3) we introduce a novel soft anchoring mechanism that fuses the source object appearances and the target scene layouts during the denoising process, (4) we show that DDPM self-attention bucketing suffices for real image editing, and finally (5) we develop automatic metrics to the task of object dragging and use them to evaluate our method quantitatively, in addition to a user study, to demonstrate its effectiveness.
2. Related Work
Localized text-to-image models.
Recently, text-to-image diffusion models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Song et al., 2020; Ramesh et al., 2022; Rombach et al., 2021; Yu et al., 2022) became a foundational tool for creative tasks (Avrahami et al., 2023d; Richardson et al., 2023; Molad et al., 2023; Frenkel et al., 2024). To add spatial control to existing text-to-image models, some works suggested training a designated localization component to take in visual layouts (Li et al., 2023; Avrahami et al., 2023c; Yang et al., 2023; Zheng et al., 2023; Zhang et al., 2023; Nie et al., 2024) while others offer training-free methods that incorporate the spatial conditioning into the diffusion sampling process (Feng et al., 2022; Chefer et al., 2023; Chen et al., 2023b; Phung et al., 2023; Bar-Tal et al., 2023). In this work, we utilize BlobGEN (Nie et al., 2024) as our base model since it has shown better spatial understanding and generation quality.
Text-to-image editing.
Soon after the emergence of text-to-image diffusion models, a plethora of methods were offered for various image editing tasks (Meng et al., 2021; Avrahami et al., 2022, 2023b; Mokady et al., 2023; Tumanyan et al., 2023; Hertz et al., 2023; Kawar et al., 2023; Cao et al., 2023; Patashnik et al., 2023; Sheynin et al., 2023). However, most of these editing methods are spatially preserving (i.e., changing the object attributes and categories), and suffer from the editing tasks requiring spatial reasoning, such as object dragging (Brooks et al., 2023; Hertz et al., 2022). Localized text-to-image models, such as GLIGEN (Li et al., 2023) and BlobGEN (Nie et al., 2024), has the potential to solve the object dragging task, but their performance is far from satisfactory without specialized designs.
Keypoint dragging.
A similar task is keypoint dragging, where users provide source and target keypoints in the image, and move the source keypoints to the target ones. For example, UserControllableLT (Endo, 2022), GANWarping (Wang et al., 2022) and DragGAN (Pan et al., 2023) employ StyleGAN (Karras et al., 2019, 2020, 2021) for editing generated images. But they work only on the narrow domain the GAN (Goodfellow et al., 2014) was trained on (e.g., human faces, churches (Yu et al., 2015)). DragDiffusion (Shi et al., 2023) propose a LoRA-based (Hu et al., 2021) method that finetunes a diffusion model given a test image and optimizes the latent noises at inference time. In contrast, our method is training-free.
Object dragging.
Different from keypoints dragging that warps the image to match the target keypoints, object dragging moves the entire object seamlessly to a new position. Object dragging was initially introduced by  (Epstein et al., 2022; Wang et al., 2021) for single-domain images generated by GANs. Diffusion self-guidance (Epstein et al., 2023) proposed to use the guidance from internal representations of a diffusion model for various editing tasks, including object dragging. DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024) developed a new classifier guidance (Dhariwal and Nichol, 2021) specifically designed for object dragging. Most of them use a general diffusion model as the base model, but our method harnesses the spatial understanding of a localized diffusion model to better tackle the object dragging task.
Object insertion.
Many works use multiple images (Ruiz et al., 2023; Gal et al., 2022; Arar et al., 2024; Alaluf et al., 2023; Voynov et al., 2023) or a single image (Avrahami et al., 2023a; Gal et al., 2023; Arar et al., 2023) of the same object for image personalization. They are also effective in tackling the task of referenced-based object insertion, in which a reference object is being inserted to a target image. AnyDoor (Chen et al., 2023a) and PaintByExample (Yang et al., 2022) train a designated encoder for this task, which can be used for object dragging by utilizing an inpainting method, as explained in Section 5. The concurrent work ObjectDrop (Winter et al., 2024) collected a high-quality tailored dataset to train a model for object removal, insertion, and dragging. Our method, however, is training-free with a pre-trained localized diffusion model.
3. Preliminaries
Figure 3. BlobGEN Architecture. BlobGEN incorporates the additional blob information into the Stable Diffusion model by adding two new layers in each attention block: masked cross-attention and gated cross-attention.
Existing large text-to-image diffusion models suffer from the prompt-following issue, making it challenging to control the visual layouts of their generation via complex prompts only. Thus, incorporating the visual layout information into these large text-to-image diffusion models can enable better object-level controllability (Li et al., 2023; Yang et al., 2023; Avrahami et al., 2023a). Among them, visual layouts are usually represented by bounding boxes (along with object categories).
More recently, BlobGEN (Nie et al., 2024) has introduced a new type of visual layouts called blob representations to guide the image synthesis, which shows more fine-grained controllability than all previous approaches. Specifically, the blob representations denote the object-level visual primitives in a scene, each of which consists of two components: blob parameters $\tau$ and blob description $S$. A blob parameter depicts a tilted ellipse using a vector of five variables $\tau=[c_{x},c_{y},a,b,\theta]$ to specify the object’s position, size and orientation, where ($c_{x}$, $c_{y}$) is the center point of the ellipse, $a$ and $b$ are the radii of its semi-major and semi-minor axes, and $\theta\in(-\pi,\pi]$ is the orientation angle of the ellipse. A blob description $S$ captures the object’s visual appearance using a region-level synthetic caption extracted by an image captioning model. Compared with bounding boxes and object categories, the blob representations can retain more detailed spatial and appearance information about the objects in a complex scene.
To incorporate blob representations into the existing Stable Diffusion model, BlobGEN adopts a similar architecture design idea to GLIGEN (Li et al., 2023) that introduces new attention layers in a gated way. To retain the prior knowledge of pre-trained models for synthesizing high-quality images, it freezes the weights of the pre-trained diffusion model and only trains the newly added layers. As demonstrated in Figure 3, BlobGEN keeps the gated self-attention module originally developed by GLIGEN while also introducing a new masked cross-attention module in each attention block. These two new layers fuse blob inputs into the model differently: In the gated self-attention layer, the blob embeddings are first passed to a pooling layer and then concatenated with the visual features, while, in the masked cross-attention layer, each blob embedding only attends to visual features in its local region as the feature maps are masked by the (rescaled) blob ellipses.
With this masking design, each blob representation and its local visual feature are trained to align with each other, and thus the model becomes more modular and disentangled. BlobGEN has demonstrated more fine-grained control over its generation. Therefore, we use BlobGEN as our network backbone for solving the object dragging task.
4. Method
Figure 4. Method Overview. Given an input image $I$, we start by extracting the blob parameters $P_{s}$ of its layout; then, by changing its layout based on the user provided target location, we get the new blob parameters $P_{d}$. By conditioning the localized text-to-image model on the respective blob representations, we iteratively denoise the source and target images ($z_{s}$ and $z_{d}$) while incorporating gated self-attention masking (Section 4.1) and soft attention anchoring (Section 4.2) in each self-attention block until we get the desired editing result $I^{\prime}$.
Our goal is to offer a solution to the problem of object dragging. To this end, we propose to leverage the spatial knowledge of blob-based text-to-image model BlobGEN (Nie et al., 2024). In Section 4.1 we start by investigating the disentanglement offered by this model. We discover significant lingering entanglement, and trace it to the gated-self attention of GLIGEN-style models. Hence, we offer an inference-time mask-based solution to this problem. In Section 4.2 we present our solution for object dragging in generated images: (1) we first utilize self-attention sharing (Cao et al., 2023; Wu et al., 2022; Geyer et al., 2023; Tewel et al., 2024) to increase the consistency of the dragged object, and (2) we propose a soft anchoring technique to improve the consistency of results. Finally, in Section 4.3 we extend our solution to real images by relying on the proposed DDPM self-attention bucketing instead of standard DDIM inversion. Our method is summarized in Figure 4.
Formally, given an input image $I$ with an object located in $(c_{x},c_{y})$ that the user wants to drag, and a desired target location $(c^{\prime}_{x},c^{\prime}_{y})$, the task of object dragging aims at moving the object to the target location while the rest of the image is left intact, up to desired environment changes (e.g., reflections) in the edited image $I^{\prime}$.
4.1. Gated Self-Attention Entanglement
Figure 5. Gated Self-Attention Leakage. Given scene descriptions of two blobs: “a photo of a rabbit” and “a photo of a cat”, the standard BlobGEN model (first row) leaks the rabbit information also to the cat blob (the first row third column), while our masked version of the gated self-attention (second row) is able to disentangle the blobs (the second row third column). In addition, we can see that the gated self-attention (second column) behaves de facto as a cross-attention layer, as the vast majority of the attention is between the text tokens $T$ and the visual tokens $V$.
As explained in Section 3, BlobGEN was trained to take a set of input blobs ${B_{1},...B_{n}}$ with corresponding text descriptions ${S_{1},...S_{n}}$ and blob parameters ${\tau_{1},...\tau_{n}}$, and generate a scene. This scene is expected to be created in a disentangled manner, i.e., the text description $S_{i}$ should correspond only to the local region depicted by $\tau_{i}$. To this end, the authors introduced a masked cross-attention layer. However, a simple investigation reveals that the generated result is not fully disentangled in practice. For example, as can be seen in Figure 5 (first row), the rabbit text description from one blob spills over to the spatial region of the cat blob.
We hypothesize that the gated self-attention modules that BlobGEN derives from GLIGEN is the root cause of entanglement. In gated self-attention, a projection layer first converts the CLIP (Radford et al., 2021) text embeddings of the text description $S_{i}$ to the text tokens $T=\{t_{1},...t_{n}\}$. They are then merged with the visual tokens $V=\{v_{1},...v_{k}\}$ into a unified set $V\cup T=\{v_{1},...v_{k},t_{1},...t_{n}\}$, which altogether are used to calculate the self-attention features, using the standard self-attention mechanism (plus a gated skip connection).
This design choice adds no constraint over the attended areas, i.e., the projected text tokens $T$ can attend to themselves and all the visual tokens $V$. To visualize this phenomenon, we average the gated self-attention maps over the diffusion process. An example is shown in Figure 5 (the second column of the first row). This visualization reveals an interesting aspect: the vast majority of attention weights is between the projected text tokens $T$ and the visual tokens $V$, and not within these sets themselves. It means that the gated self-attention layer behaves as a de facto cross-attention layer.
We examine the attention between the projected text token $t_{i}$ and all the visual tokens $V$. This is a $K$-dimensional vector, which we first reshape into two dimensions $\sqrt{K}\times\sqrt{K}$, and then resize to a canonical size. We term these maps “reshaped self-attention”, which are averaged over all the denoising steps. This visualization, as shown in Figure 5 (last two columns of the first row), reveals that text tokens indeed attend to undesired areas: the “rabbit” text token attends to the visual features in both the “rabbit” and “cat” blob regions, leading to an entangled generation. For more details about the visualizations, please refer to the supplementary material.
To this end, we suggest an inference-time solution to the entanglement problem: given $n$ different input blobs with the corresponding parameters ${\tau_{1},...\tau_{n}}$ we first convert them into $n$ masks ${M_{1},...M_{n}}$ of $512\times 512$ resolution. Then, during the diffusion process, for each self-attention layer and for each projected text token $t_{i}$, we reshape the mask $M_{i}$ to the corresponding spatial size of the layer, and use it to mask the area of the gated self-attention between the projected text token $t_{i}$ and the visual tokens $V$. This way, we can prevent the token $t_{i}$ from attending to undesired areas at the inference time.
4.2. Consistent Object Dragging for Generated Images
Figure 6. Self-Attention Soft Anchoring. Given the source blob $B_{s}$ and target blob $B_{d}$, we start by extracting the self-attention outputs $O_{s}$ and $O_{d}$ correspondingly, then, during the first $\rho$ iterations, we blend these maps according to the timestep ratio $f=\frac{t}{T}$ where $t$ is the current timestep and $T$ is the total number of timesteps. Then, after the anchor map $O_{a}$ is calculated, we use it for determining the position of the new blob, while taking the appearance from the corresponding $O_{s}$ map using nearest-neighbor copying.
Now, we first focus on tackling the object dragging problem for generated images from the localized model: given a scene represented by $n$ blob inputs ${B_{1},...B_{n}}$, we change the parameters $\tau_{s}$ of one blob $B_{s}$ to $\tau_{d}$ with a different spatial location such that the $s^{\text{th}}$ object in the generated image will be relocated to the designated location, without changing the appearance of all other objects and the background (barring direct interactions with the object, e.g., shadows).
To preserve the high-level object appearance, we adopt the self-attention sharing mechanism (Cao et al., 2023; Wu et al., 2022): we iteratively generate the source image using the source parameters $\tau_{s}$ in parallel to the target image with the $\tau_{d}$ parameters. Then, we replace the self-attention keys $K_{d}$ and values $V_{d}$ from the target image in each self-attention layer and each denoising step by the keys and values $K_{s},V_{s}$ from the source image.
However, this mechanism alone does not fully preserve the fine-grained details of the source image, so we propose adding a novel soft anchoring mechanism: the motivation is that the generated source image already contains the information needed for generating the target image, we can take advantage of the self-attention layers output (i.e., attention features) in the local region that corresponds to the source blob. The soft anchoring is designed to fuse the object appearance information represented by the attention features within the source blob and the positional information indicated by the target blob. Specifically, in the first $\rho$ steps of the denoising process, we perform an adaptive, soft blending of the attention features of the generated target image with the features of the source image. The interpolation coefficient is time-dependent: we take more visual appearance from the source image in the beginning but more spatial information from the target image in the later steps, as depicted in Figure 6. Formally, for each denoising step $t\in[T,T-1,...,T-\rho+1]$ and for each self-attention layer, the interpolated self-attention output of the target image is:
$O_{a}=O_{s}*f+O_{d}*(1-f);f=\frac{t}{T}$
where $O_{s}$ is the self-attention output of the generated source image, $O_{d}$ is the self-attention output of the generated target image, and $T$ is the total number of denoising steps. The length of soft blending is controlled by the hyperparameter $\rho$.
Next, during the last $T-\rho$ steps of the denoising process, we use the soft blending result $O_{a}$ as anchor points for the target object. In each denoising step $t\in[T-\rho,...,2,1]$ and each self-attention layer, we perform the nearest-neighbor copying: each entry from the anchor attention features $O_{a}$ within the target blob $B_{d}$ is replaced by its nearest-neighbor entry from the source attention features $O_{s}$ within the source blob $B_{s}$. The nearest-neighbor entry is obtained by measuring the normalized cosine similarity. Formally,
$(O_{a})_{(j,k)\in B_{d}}=(O_{s})_{\textit{NN}(j,k)\in B_{s}}$
where ${(j,k)\in B_{d}}$ represents the set of coordinates for each entry from $O_{a}$ within the target blob $B_{d}$ and $\textit{NN}(j,k)\in B_{s}$ denotes the set of coordinates for each nearest-neighbor entry from $O_{s}$ within the source blob $B_{s}$.
4.3. Extension for Real Images
In order to extend our method for dragging objects in real images, we first extract the blobs parameters as we explain later, then, we need to invert the image. However, we found that directly applying the DDIM inversion (Song et al., 2020) in a localized model is not able to preserve the details of the input image, even without classifier-free guidance (Ho, 2022).
Recall that when dealing with generated images, the input signal from the source image is fed in our pipeline through its self-attention outputs. Hence, we only need to extract the self-attention features in different attention layers and different denoising steps from the real image, rather than an actual inversion that searches for the optimal latent noises. To this end, we propose the DDPM self-attention bucketing: we first add independent noises with various scales to the real image, where the noise scale corresponds to a time step in the DDPM forward process. The noisy images at every time step, along with the above extracted blobs, are then passed to the localized model to get self-attention outputs in every attention layer, as needed. Note that the DDPM self-attention bucketing is specifically designed for the object dragging task, where we aim to preserve the visual details of the real image. It may not be suitable for other image editing tasks that change the object appearance or category.
For extracting the blobs representations from real images, we utilize ODISE (Xu et al., 2023) to get instance segmentation maps, then we use an ellipse fitting optimization with the goal of maximizing the Intersection Over Union (IOU) between the ellipse and the generated mask. Finally, we crop a local region around each blob and use LLaVA-1.5 (Liu et al., 2023) for the local captioning.
Finally, in order to better preserve the background, we incorporated the Blended Latent Diffusion (Avrahami et al., 2022, 2023b) method into our process in which the background pixels are being integrated into the diffusion process in order to seamlessly blend the generated result in the original scene. For more details, please refer to the supplementary material.
5. Experiments
In Section 5.1, we compare our method against several baselines, both qualitatively and quantitatively. Next, in Section 5.2 we describe the user study on various methods and present the outcome. Lastly, in Section 5.3, we show the ablation study results to highlight the importance of each component.
5.1. Qualitative and Quantitative Comparison
Table 1. Quantitative Comparison. We compare our method against the baselines in terms of foreground similarity (higher is better), object traces (lower is better) and realism (lower is better). As can be seen, DiffEditor (Mou et al., 2024) and DragonDiffusion (Mou et al., 2023) struggle with object traces as they suffer from the object traces issue. PBE (Yang et al., 2022), Anydoor (Chen et al., 2023a), DragDiffusion (Shi et al., 2023) and Diffusion SG (Epstein et al., 2023) struggle with foreground similarity as they tend not to drag the object. In contrast, our method significantly outperforms all the baselines in terms of object traces and also achieves higher foreground similarity with comparable image realism.
Method Foreground ()($\uparrow$)( ↑ ) Traces ()($\downarrow$)( ↓ ) Realism ()($\downarrow$)( ↓ ) PBE (Yang et al., 2022) 0.614 0.446 0.0029 DiffusionSG (Epstein et al., 2023) 0.515 0.459 0.0096 Anydoor (Chen et al., 2023a) 0.684 0.454 0.0035 DragDiffusion (Shi et al., 2023) 0.738 0.603 0.0011 DragonDiffusion (Mou et al., 2023) 0.818 0.773 0.0009 DiffEditor (Mou et al., 2024) 0.826 0.604 0.0008 DiffUHaul (ours) 0.835 0.32 0.0008
Table 2. Ablation Study. We ablate the following components of our method: (1) w/o gated self-attention (GSA) masking, (2) w/o self-attention (SA) sharing, (3) w/o soft attention anchoring and (4) w/o DDPM noising. As can be seen, removing the (1) GSA masking harms the foreground similarity, as leakages from neighboring blobs can interfere. Removing the (2) SA sharing or the (3) soft attention anchoring harms the foreground similarity as well, as it reduces the similarity the input image. Removing the DDPM SA bucketing slightly improves the object traces but significantly harms the foreground similarity, as the details of source images are not well preserved.
Method Foreground ()($\uparrow$)( ↑ ) Traces ()($\downarrow$)( ↓ ) Realism ()($\downarrow$)( ↓ ) DiffUHaul (ours) 0.835 0.320 0.0008 w/o GSA masking 0.823 0.320 0.0008 w/o SA sharing 0.755 0.314 0.0014 w/o soft attention anchoring 0.780 0.322 0.0008 w/o DDPM SA bucketing 0.675 0.298 0.0029
We compared our method against the most relevant available object dragging baselines. Paint-By-Example (PBE) (Yang et al., 2022) and AnyDoor (Chen et al., 2023a) present a way of adding an object to an image. To use them for object dragging, we crop the object from the source image, apply the image inpaintng in the cropped region, and then add the object in the new location. Diffusion Self-Guidance (Diffusion SG) (Epstein et al., 2023) tackles the general image editing tasks via attention guidance, which can be tailored to object dragging. DragDiffusion (Shi et al., 2023) is designed for the task of keypoint-based dragging, which we can convert into object dragging by selecting multiple points on the source object. Finally, DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024) directly tackle the problem of object dragging. For more details, please see the supplementary material.
As can be seen in Figure 8, PBE (Yang et al., 2022), Anydoor (Chen et al., 2023a) and DiffusionSG (Epstein et al., 2023) can hardly preserve the appearance of the edited object and always have undesirable objects or artifacts left in the source location, indicating that existing general-purpose image editing methods tend to completely fail in the object dragging task. DragDiffusion (Shi et al., 2023) struggles with moving the object to the target location, while DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024) often suffers from the object traces issue, where the object appears in both source and target locations. In contrast, our method strikes the best balance between effectively dragging the object to the right position and preserving its visual appearance.
To quantify the performance of our method and baselines, we prepare a specialized evaluation dataset based on the COCO (Lin et al., 2014) validation set. We first filter it to contain only images that have a single “thing” object with a prominent size. Then, we use the same blobs extraction pipeline as explained in Section 4.3. For the object dragging task, we randomly sample a new location in the pixel space as the center of the target blob $B_{d}$. For each sample, we calculate 8 different target drag locations, resulting in a total dataset of 6,048 samples. For more details, please read the supplementary material. In Figure 9, we provide a qualitative comparison on the automatic dataset, where we make similar observations as before.
Based on this new dataset, we propose three evaluation metrics: foreground similarity, object traces and realism. Foreground similarity quantifies whether the source object indeed dragged to the target location without appearance changes. To this end, we crop a tight box area around the source blob $B_{s}$ in the source image $I_{s}$ and around the target blob $B_{d}$ in the target image $I^{\prime}$, respectively, and pass the crops to DINOv2 (Oquab et al., 2023) to measure the perceptual similarity after aligning them to a canonical position and masking the background. We strive to maximize this metric. To measure the object traces phenomenon, we crop a tight box area around the source blob $B_{s}$ in the source image $I_{s}$ and around the source blob $B_{s}$ in the target image $I^{\prime}$. Next, we masked the target blob $B_{d}$ area in the target image $I^{\prime}$. Similarly, we utilize DINOv2 (Oquab et al., 2023) to measure the perceptual similarity between the crops. We strive to minimize this metric. Lastly, to measure the realism of the edited image, we utilized KID score (Binkowski et al., 2018) of sets of 672 real and generated images. For more details, please read the supplementary material.
As can be seen in Table 1, DiffEditor (Mou et al., 2024) and DragonDiffusion (Mou et al., 2023) rank high in object traces as they suffer from object traces problem. PBE (Yang et al., 2022), Anydoor (Chen et al., 2023a), DragDiffusion (Shi et al., 2023) and Diffusion SG (Epstein et al., 2023) struggle with foreground similarity as they tend not to drag the object. On the other hand, our method significantly outperforms all the baselines in terms of object traces, which demonstrates the robustness of our method. In addition, it achieves higher foreground similarity and is on par in terms of image realism. These results are supported by the qualitative comparison.
5.2. User Study
Table 3. User Study. We compare our method against the baselines using the standard two-alternative forced-choice format. Users were asked to rate which editing result is better (Ours vs. the baseline) in terms of: (1) dragging the object to the desired location (2) leaving no traces of the original object, (3) realism and (4) overall edit quality. The number represents the win rate of our method over each of the baselines. As we can see, our method wins the baselines in all terms more than the random win rate of 50%.
Ours vs Dragging ()($\uparrow$)( ↑ ) No traces ()($\uparrow$)( ↑ ) Realism ()($\uparrow$)( ↑ ) Overall ()($\uparrow$)( ↑ ) PBE (Yang et al., 2022) 82.14% 78.57% 82.58% 79.68% DiffusionSG (Epstein et al., 2023) 79.46% 76.78% 77.45% 77.45% Anydoor (Chen et al., 2023a) 81.02% 76.78% 81.91% 81.25% DragDiffusion (Shi et al., 2023) 64.73% 58.92% 59.82% 61.16% DragonDiffusion (Mou et al., 2023) 77.00% 75.22% 81.02% 76.56% DiffEditor (Mou et al., 2024) 73.43% 70.08% 70.08% 76.33%
We conduct an extensive user study using the Amazon Mechanical Turk (AMT) platform (Amazon, 2024), where the test examples are also sampled from the automatically extracted dataset as explained in Section 5.1. We compare all the baselines using the standard two-alternative forced-choice format. Users were given the source image, the edit instructions and two edited images: one from our method and another one from a baseline. For each comparison, users were asked to rate which edited image is better in terms of: (1) dragging the object to the desired location (2) leaving no traces of the original object, (3) realism and (4) overall edit quality (i.e., taking all the aspects into account). As can be seen in Table 3, our method is preferred over all the baselines in terms of the overall edit quality and different individual perspectives. This observation aligns well with our automatic metrics. The user study suggests that DragDiffusion is the second-strongest baseline, it may be due to the fact that it also result with realistic images, as it also avoids leaving traces of the dragged object, which the automatically calculated KID do not take into account. For more details and statistical significance analysis, please read the supplementary material.
5.3. Ablation Study
We perform the ablation study for the following components of our method: (1) Without gated self-attention masking — we remove the gated self-attention masking that is described in Section 4.1. (2) Without self-attention sharing — we remove the self-attention sharing component. (3) Without soft attention anchoring — we remove the soft attention anchoring that is described in Section 4.2. (4) Without DDPM noising — we replace the DDPM noising that is described in Section 4.3 with a DDIM inversion (Song et al., 2020).
We use the same automatic evaluation metrics as described in Section 5.1 to quantify the importance of each component. As can be seen in Table 2, removing the (1) GSA masking harms the foreground similarity, as leakages from neighboring blobs can interfere the visual appearances of the focused object. Removing the (2) SA sharing or the (3) soft attention anchoring harms the foreground similarity as well, as it reduces the similarity to the input image. Removing the DDPM noising slightly improves the object traces, but it significantly harms the foreground similarity, as the reconstructed image itself has changed significantly. For a qualitative visualization of the ablation study, please refer to the supplementary material.
6. Limitations and Conclusions
(a) Rotation
Refer to caption Refer to caption Refer to caption Refer to caption
Original Rotation 1 Rotation 2 Rotation 3
(b) Resize
Refer to caption Refer to caption Refer to caption Refer to caption
Original Resize 1 Resize 2 Resize 3
(c) Collision
Refer to caption Refer to caption Refer to caption Refer to caption
Original Drag 1 Drag 2 Drag 3
Figure 7. Limitations. Our method suffers from the following limitations: (a) We found our method to be incapable of rotating objects, and instead stretch the object to fit the new blob shape without changing the orientation. (b) We found our method to struggle with resizing object, especially in large resizes (e.g., Resize 3 in the second row). (c) We found our method to struggle to handle colliding object while dragging, which may result with a hybrid between the objects (e.g., Drag 2 in the third row) or one object being merged (e.g., Drag 3 in the third row).
Our method suffers from the following limitations that are depicted in Figure 7: (a) We found our diffusion anchoring technique introduced in Section 4.2 to be incapable of rotating objects, and instead, as can be seen in Figure 7(a), stretch the object to fit the new blob shape without changing the orientation, this may be caused due to the fact that rotation involves understanding the 3D structure, which is not reflected by the self-attention nearest-neighbor copying. (b) We found our method to struggle with resizing object, especially in large resizes, as can be seen in Figure 7(b) Resize 3. (c) We found our method to struggle to handle colliding object while dragging, which may result with a hybrid between the objects (Figure 7(c) Drag 2) or one object being merged (Figure 7(c) Drag 3).
In conclusion, we presented DiffUHall, our solution to the seemingly straightforward task of object dragging. We demonstrated that the spatial understanding of the localized BlobGEN can be harnessed to this task, using our novel diffusion anchoring technique that manages to merge the location signal from the model with the object appearance signal from the input image. We hope that our contributions will serve as a valuable tools not only for object dragging but also for other creative tasks in the future.
Input
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PBE
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Diffusion SG
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
AnyDoor
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DragDiffusion
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DragonDiffusion
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DiffEditor
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8. Qualitative Comparison. We compared our method against several baselines on both generated (first three columns) and real images (second three columns). The source and target locations are denoted by red and green points, respectively. As can be seen, PBE (Yang et al., 2022), DiffusionSG (Epstein et al., 2023) and Anydoor (Chen et al., 2023a) mainly suffer from a bad preservation of the foreground object. DragDiffusion (Shi et al., 2023) struggles with dragging the object, while DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024) suffers from object traces. Our method, on the other hand, strikes the balance between dragging the object and preserving its identity.
Input
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PBE
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Diffusion SG
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
AnyDoor
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DragDiffusion
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DragonDiffusion
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DiffEditor
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9. Qualitative Automatic Comparison. As explained in Section 5.1, we used a filtered version of COCO validation set (Lin et al., 2014). The source and target locations are denoted by red and green points, respectively. As can be seen, PBE (Yang et al., 2022), DiffusionSG (Epstein et al., 2023) and Anydoor (Chen et al., 2023a) mainly suffer from a bad preservation of the foreground object. DragDiffusion (Shi et al., 2023) struggles with dragging the object, while DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024) suffers from object traces. Our method, on the other hand, strikes the balance between dragging the object and preserving its identity.
Appendix A Implementation Details
In Section A.1 we start by providing the implementation details of our method. Next, in Section A.2 we provide the baselines’ implementation details. Later, in Section A.3 we provide the implementation details of the automatic metrics we used. Finally, in Section A.4 we provide the detail of the user study we conducted.
A.1. Implementation Details of Our Method
Below, we provide the full implementation details of our method: in Section A.1.1 we start by providing the implementation details for the gated self-attention visualization we used, next, in Section A.1.2 we provide the implementation details of the soft self-attention anchoring we used, and finally, in Section A.1.3 we explain about the Blended Latent Diffusion integration.
A.1.1. Gated Self-Attention Visualization Implementation Details
Figure 10. Full Gated Self-Attention Visualization. We provide our visualization for the gated self-attention layers, before our inference time masking (left) and after it (right). As can be seen, even though there is no constraint on the attention distribution (as in any other self-attention layer), we found empirically that the vast majority of the attention is being formed between the used pooled textual tokens (the first part of $T$) and the visual tokens $V$. And not between the sets themselves. This behavior suggests that this kind of layers behaves de facto as cross attention layers. Our inference-time masking mechanism (right) constrain the attention between the textual tokens and only their corresponding visual tokens within their blobs.
As explained in Section 4.1, we offer a method to visualize the gated self-attention layer of GLIGEN (Li et al., 2023) and BlobGEN (Nie et al., 2024). In gated self-attention, a projection layer first converts the CLIP (Radford et al., 2021) text embeddings of the text description $S_{i}$ to the text tokens $T=\{t_{1},...t_{n}\}$. They are then merged with the visual tokens $V=\{v_{1},...v_{k}\}$ into a unified set $V\cup T=\{v_{1},...v_{k},t_{1},...t_{n}\}$, which altogether are used to calculate the self-attention features, using the standard self-attention mechanism (plus a gated skip connection).
We examine the attention between the projected text token $t_{i}$ and all the visual tokens $V$. This is a $K$-dimensional vector, which we first reshape into two dimensions $\sqrt{K}\times\sqrt{K}$, and then resize to a canonical size. We term these maps “reshaped self-attention”, which are averaged over all the denoising steps. We average these maps across all the diffusion steps and all the layers, by resizing them to a canonical size.
In Figure 10 we provide a full visualization of these maps. As can be seen, even though there is no constraint on the attention distribution (as in any other self-attention layer), we found empirically that the vast majority of the attention is being formed between the used pooled textual tokens (the first part of $T$) and the visual tokens $V$. There is little interaction within the individual sets themselves. This behavior suggests that this kind of self-attention layers behaves as de facto cross attention layers. Our inference-time masking mechanism (right) further constrains each textual token to only attend to its corresponding visual token within the blob region.
Please note that these kind of visualizations and the masking manipulations are different from the those common in the text-to-image diffusion-based models literature (Hertz et al., 2022; Avrahami et al., 2023a; Chefer et al., 2023) as we do not manipulate the traditional cross-attention layers but the gated self-attention layers.
A.1.2. Soft Self-Attention Anchoring Implementation Details
As explained in Section 4.2, we propose the soft self-attention anchoring to fuse the spatial information from the localized model and the appearance information from the input source image. Specifically, in the first $\rho=\frac{T}{2}$ steps of the denoising process, we perform an adaptive, soft blending of the attention features of the generated target image with the features of the source image. The interpolation coefficient is time-dependent: we take more visual appearance from the source image in the beginning but more spatial information from the target image in the later steps.
Next, during the last $T-\rho$ steps of the denoising process, we use the soft blending result $O_{a}$ as anchor points for the target object. In each denoising step $t\in[T-\rho,...,2,1]$ and each self-attention layer, we perform the nearest-neighbor copying: each entry from the anchor attention features $O_{a}$ within the target blob $B_{d}$ is replaced by its nearest-neighbor entry from the source attention features $O_{s}$ within the source blob $B_{s}$. To calculate the nearest-neighbor, we normalize each self-attention entry of $O$ and calculate its cosine similarity with each entry of $O_{s}$. Please note that the nearest-neighbor operation is only within the source blobs $B_{s}$ and destination blob $B_{d}$ boundaries. To calculate the these blobs, we reshape them to the corresponding self-attention size of each layer.
A.1.3. Blended Latent Diffusion Integration
Blended Latent Diffusion (Avrahami et al., 2022, 2023b) is a method designed for localized image editing using text-to-image diffusion models. the input image is fused into the diffusion process along with an input mask to preserve it background, while encouraging the generated content (in the unmasked area) to be consistent to the background. We also use this method in our pipeline of editing real images, as introduced in Section 4.3. Given the source blob $B_{s}$ and the destination blob $B_{d}$ provided by the user, we take the union blob that contains both of them $B_{u}=B_{s}\cup B_{d}$, and morphologically dilate it with a kernel of a size of $50\times 50$. We treat this dilated blob as the editable area, which we provide to the Blended Latent Diffusion method to edit real images during the entire diffusion process (i.e.  the hyperparameter of noising diffusion steps $k=T$, where $T$ is the total number of diffusion steps).
A.2. Implementation Details of Baselines
Table 4. Inference Time Comparison. We report the inference time of the baselines and our method of editing a single $512\times 512$ image. All the reported running times we calculated using a single NVIDIA A100 GPU.
Method Inference time $(sec)$
PBE (Yang et al., 2022) 9 sec
Diffusion SG (Epstein et al., 2023) 14 sec
Anydoor (Chen et al., 2023a) 9 sec
DragDiffusion (Shi et al., 2023) 148 sec
DragonDiffusion (Mou et al., 2023) 13 sec
DiffEditor (Mou et al., 2024) 13 sec
DiffUHaul (ours) 13 sec
As described in Section 5.1, we compare our method against the available, most relevant object dragging baselines: Paint-By-Example (Yang et al., 2022), AnyDoor (Chen et al., 2023a), Diffusion Self-Guidance (Epstein et al., 2023), DragDiffusion (Shi et al., 2023), DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024). Out of these, Diffusion Self-Guidance, DragonDiffusion and DiffEditor directly support the task of object dragging. The rest of these baselines need some adaptations to our problem, as described below.
Paint-By-Example and AnyDoor present a way to add an object to an image. Hence, in order to convert it to our problem setting we constructed a designated pipline: we started taking the source image $I$ and inpaint the source area blob $B_{s}$ using Stable Diffusion Inpaint (von Platen et al., 2022), to get an inpainted version $\hat{I}$. Then, we used Paint-By-Example/Anydoor to inpaint the new image $\hat{I}$ again, in the target blob area $B_{d}$ by providing the original object in the original image $I$ as a reference.
DragDiffusion is originally designed to tackle the problem of keypoint-based dragging. Thus, in order to adapt it to our method, we take the centroid of the source blob $B_{s}$ as well as other points sampled inside the source blob region, and then translate them to the target blob $B_{d}$.
We used the official baseline implementations with a comparable backbone of Stable Diffusion v1 (Rombach et al., 2021) using 50 DDIM diffusion steps, except Diffusion Self-Guidance (Epstein et al., 2022), of which the only available implementation is based upon SDXL (Podell et al., 2023).
In Table 4 we report the inference time of our method and the baselines for a single image editing using an NVIDIA A100 GPU. All the methods takes around 10 seconds, except DragDiffusion that is using an extensive LoRA (Hu et al., 2021; Horwitz et al., 2024) training and latent optimization, which increases the inference time significantly.
We used the following third-party packages in this research:
A.3. Implementation Details of Automatic Metrics
As described in Section 5.1, in order to automatically compare our method against baselines quantitatively, we utilized COCO (Lin et al., 2014) validation dataset. We filtered it to contain only images with a main “thing” class object (the number of “stuff” object is unbounded) that occupies at least 5% of the image size, but not more than 25% of the image size. Then, we utilize ODISE (Xu et al., 2023) to get instance segmentation maps, then we use an ellipse fitting optimization with the goal of maximizing the Intersection Over Union (IOU) between the ellipse and the generated mask. Next, we crop a local region around each blob and use LLaVA-1.5 (Liu et al., 2023) for the local captioning. Finally, we choose a random location for the target blob $B_{d}$ that is at least 64 pixels. This resulted with 672 filtered images and 8 target blob locations per image, which is a total of 6,048 evaluated samples per baseline.
Next, we propose using three metrics: foreground similarity, object traces and realism. Foreground similarity quantifies whether the source object is indeed dragged to the target location. To this end, we crop a tight square area around the source blob $B_{s}$ in the source image $I$, and the target blob $B_{d}$ in the target image $T_{t}$. Next, we align the crops to a canonical position and mask the background in these crops by aligning the object to the left side of the image (in order to avoid translation artifacts). Finally, we utilize DINOv2 (Oquab et al., 2023) to measure the perceptual similarity between these crops. We strive to maximize this metric.
Similarly, in order to measure the object duplication phenomenon, we crop a tight square area around the source blob $B_{s}$ in the source image $I$, and around the source blob $B_{s}$ in the target image $I^{\prime}$. Next, we mask the target blob $B_{d}$ area in the target image $I^{\prime}$. Finally, we utilize again DINOv2 (Oquab et al., 2023) to measure the perceptual similarity between these crops. We strive to minimize this metric. Lastly, in order to measure the realism of the image, we compute the KID score (Binkowski et al., 2018) using 672 real and generated images. The reason of using KID instead of the FID score (Heusel et al., 2017) is that it better aligns with human perception of image generation quality when the provided real and fake sets are small.
Figure 11. User Study Trial. We provide an example of one trial task in the user study we conducted using Amazon Mechanical Turk (AMT) (Amazon, 2024). The users were asked four questions of a two-alternative forced-choice format. The full instructions can be seen in Figure 12.
Figure 12. User Study Instructions. We provide the full instructions for the user study we conducted using Amazon Mechanical Turk (AMT) (Amazon, 2024), to compare our method with each baseline.
A.4. User Study Implementation Details
Table 5. User Study Statistical Significance. A binomial statistical test of the user study results suggests that our results are statistically significant (p-value ¡ 5%)
Ours vs Dragging ()($\uparrow$)( ↑ ) No traces ()($\uparrow$)( ↑ ) Realism ()($\uparrow$)( ↑ ) Overall ()($\uparrow$)( ↑ ) p-value p-value p-value p-value PBE (Yang et al., 2022) ¡ 1e-8 ¡ 1e-8 ¡ 1e-8 ¡ 1e-8 DiffusionSG (Epstein et al., 2023) ¡ 1e-8 ¡ 1e-8 ¡ 1e-8 ¡ 1e-8 Anydoor (Chen et al., 2023a) ¡ 1e-8 ¡ 1e-8 ¡ 1e-8 ¡ 1e-8 DragDiffusion (Shi et al., 2023) ¡ 1e-8 ¡ 2e-5 ¡ 6e-6 ¡ 5e-7 DragonDiffusion (Mou et al., 2023) ¡ 1e-8 ¡ 1e-8 ¡ 1e-8 ¡ 1e-8 DiffEditor (Mou et al., 2024) ¡ 1e-8 ¡ 1e-8 ¡ 1e-8 ¡ 1e-8
As explained in Section 5.2 We conduct an extensive user study using the Amazon Mechanical Turk (AMT) platform (Amazon, 2024). We use the automatically extracted dataset as explained in Section 5.1 in the main paper. We compare all the baselines using the standard two-alternative forced-choice format. Users are instructed the following “In the following image, we are interested in moving the {CATEGORY} from the original location, indicated by the red dot, to the target location, indicated by a green dot.” where {CATEGORY} is the  (Lin et al., 2014) object class category. Then, the users are given two editing results: our method and one of the baselines, and are asked: (1) “Which of the results is better in moving the {CATEGORY} to the target location?”, (2) “Which of the results is better in leaving no traces of the {CATEGORY} in the original location?”, (3) “Which of the results looks more realistic?” and (4) “Which of the results is better overall?”. The users are also given detailed instructions with examples. An example of one trail can be seen in Figure 11, and the full instructions can be seen in Figure 12.
We gather 7 ratings per sample, resulting 448 ratings per baseline, totaling 2,688 responses. The time allotted per task is one hour, to allow the raters to properly evaluate the results without time pressure. A binomial statistical test of the user study results, as presented in Table 5, suggesting that our results are statistically significant (p-value ¡ 5%).
Appendix B Additional Results
Input
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PBE
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Diffusion SG
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
AnyDoor
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DragDiffusion
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DragonDiffusion
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DiffEditor
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13. Additional Qualitative Automatic Comparison. As explained in Section 5.1, we used a filtered version of COCO validation set (Lin et al., 2014). The source and target locations are denoted by red and green points, respectively. As can be seen, PBE (Yang et al., 2022), DiffusionSG (Epstein et al., 2023) and Anydoor (Chen et al., 2023a) mainly suffer from a bad preservation of the foreground object. DragDiffusion (Shi et al., 2023) struggles with dragging the object, while DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024) suffers from object traces. Our method, on the other hand, strikes the balance between dragging the object and preserving its identity.
Input
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/o GSA masking
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/o SA sharing
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/o anchoring
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/o DDPM
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14. Qualitative Results in Ablation Study. As explained in Section 5.3, we ablate four key components of our method: (a) w/o GSA masking, (b) w/o SA sharing, (c) w/o soft attention anchoring and (d) w/o DDPM SA attention. As can be seen, all these components improve the foreground object consistency.
In Figure 13 we provide an additional qualitative comparison of our method against the baselines on the automatically extracted dataset (as explained in Section A.3). As can be seen, PBE (Yang et al., 2022), DiffusionSG (Epstein et al., 2023) and Anydoor (Chen et al., 2023a) mainly suffer from a bad preservation of the foreground object. DragDiffusion (Shi et al., 2023) struggles with dragging the object, while DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024) suffers from object traces. Our method, on the other hand, strikes the balance between dragging the object and preserving its identity.
In addition, in Figure 14 we provide a qualitative visualization of the ablation study we conducted. As can be seen, PBE (Yang et al., 2022), DiffusionSG (Epstein et al., 2023) and Anydoor (Chen et al., 2023a) mainly suffer from a bad preservation of the foreground object. DragDiffusion (Shi et al., 2023) struggles with dragging the object, while DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024) suffers from object traces. Our method, on the other hand, strikes the balance between dragging the object and preserving its identity.
Appendix C Societal Impact
We believe that the advent of technology enabling seamless object dragging within images holds tremendous promise for a wide array of creative and practical uses. It may democratize content manipulation for individuals lacking expertise and artistic skills. Furthermore, we believe that this method may present an invaluable tool for professional artists by expediting their creative processes without compromising quality.
Conversely, akin to other generative AI technologies, this method is susceptible to misuse, potentially giving rise to the creation of deceptive and misleading visual content. The ease and accessibility afforded by this technology could amplify concerns regarding the authenticity and trustworthiness of visual media in various contexts, including journalism, advertising, and social media which may erode the public trust in such content. Therefore, while recognizing its transformative potential, it is imperative to remain vigilant and implement appropriate safeguards to mitigate the proliferation of misinformation and uphold ethical standards in content creation and dissemination.

### v2

@@
 Object Draggining, Image Editing
-journal: TOGsubmissionid: 0ccs: Computing methodologies Machine learningccs: Computing methodologies Computer graphics
+journalyear: 2024copyright: rightsretainedconference: SIGGRAPH Asia 2024 Conference Papers; December 3–6, 2024; Tokyo, Japanbooktitle: SIGGRAPH Asia 2024 Conference Papers (SA Conference Papers ’24), December 3–6, 2024, Tokyo, Japandoi: 10.1145/3680528.3687590isbn: 979-8-4007-1131-2/24/12submissionid: 308ccs: Computing methodologies Machine learningccs: Computing methodologies Computer graphics
 Figure 1.
@@
 Localized text-to-image models, such as GLIGEN (Li et al., 2023) and BlobGEN (Nie et al., 2024), has the potential to solve the object dragging task, but their performance is far from satisfactory without specialized designs.
+Concurrently, Diffusion Handles (Pandey et al., 2023) offers 3D object edits using a depth-to-image diffusion model and by performing manipulations on the diffusion activations in 3D.
+In addition, Magic Fixup (Alzayer et al., 2024) offers a model that given a coarsely edited image, synthesizes a photorealistic version of it, by leveraging a video dataset, this way they manage to offer a way to edit an image coarsely, and then harmonize the result.
 Keypoint dragging.
@@
 In contrast, our method is training-free.
+Concurrently, EasyDrag (Hou et al., 2024) improves DragDiffusion (Shi et al., 2023) by replacing the LoRA training with reference guidance.
 Object dragging.
@@
 Gated Self-Attention Leakage.
-Given scene descriptions of two blobs: “a photo of a rabbit” and “a photo of a cat”, the standard BlobGEN model (first row) leaks the rabbit information also to the cat blob (the first row third column), while our masked version of the gated self-attention (second row) is able to disentangle the blobs (the second row third column).
+Given scene descriptions of two blobs: “a photo of a rabbit” and “a photo of a cat”, we can see that the standard BlobGEN model (the first column in the first row) generates two rabbits instead of a cat and a rabbit, we then visualize the gated self-attention layers, as explained in Section 4.1.
+As can be seen, the standard BlobGEN model (first row) leaks the rabbit information also to the cat blob (the first row third column), while our masked version of the gated self-attention (second row) is able to disentangle the blobs (the second row third column).
 In addition, we can see that the gated self-attention (second column) behaves de facto as a cross-attention layer, as the vast majority of the attention is between the text tokens $T$ and the visual tokens $V$.
@@
 In order to extend our method for dragging objects in real images, we first extract the blobs parameters as we explain later, then, we need to invert the image.
-However, we found that directly applying the DDIM inversion (Song et al., 2020) in a localized model is not able to preserve the details of the input image, even without classifier-free guidance (Ho, 2022).
+However, we found that directly applying theg DDIM inversion (Song et al., 2020) in a localized model is not able to preserve the details of the input image, even without classifier-free guidance (Ho, 2022).
+Using more advanced inversion methods (Huberman-Spiegelglas et al., 2023; Qi et al., 2023) will not work in our case, as they preserve the general structure of the scene, whereas we are interested in changing the scene layout significantly.
 Recall that when dealing with generated images, the input signal from the source image is fed in our pipeline through its self-attention outputs.
@@
 Paint-By-Example (PBE) (Yang et al., 2022) and AnyDoor (Chen et al., 2023a) present a way of adding an object to an image.
-To use them for object dragging, we crop the object from the source image, apply the image inpaintng in the cropped region, and then add the object in the new location.
+To use them for object dragging, we crop the object from the source image, apply the image inpainting in the cropped region, and then add the object in the new location.
 Diffusion Self-Guidance (Diffusion SG) (Epstein et al., 2023) tackles the general image editing tasks via attention guidance, which can be tailored to object dragging.
@@
 To measure the object traces phenomenon, we crop a tight box area around the source blob $B_{s}$ in the source image $I_{s}$ and around the source blob $B_{s}$ in the target image $I^{\prime}$.
-Next, we masked the target blob $B_{d}$ area in the target image $I^{\prime}$.
+Next, we mask the target blob $B_{d}$ area in the target image $I^{\prime}$.
 Similarly, we utilize DINOv2 (Oquab et al., 2023) to measure the perceptual similarity between the crops.
 We strive to minimize this metric.
-Lastly, to measure the realism of the edited image, we utilized KID score (Binkowski et al., 2018) of sets of 672 real and generated images.
+Lastly, to measure the realism of the edited image, we utilize KID score (Binkowski et al., 2018) of sets of 672 real and generated images.
 For more details, please read the supplementary material.
@@
 This observation aligns well with our automatic metrics.
-The user study suggests that DragDiffusion is the second-strongest baseline, it may be due to the fact that it also result with realistic images, as it also avoids leaving traces of the dragged object, which the automatically calculated KID do not take into account.
+The user study suggests that DragDiffusion is the second-strongest baseline, it may be due to the fact that it also results with realistic images, as it also avoids leaving traces of the dragged object, which the automatically calculated KID do not take into account.
 For more details and statistical significance analysis, please read the supplementary material.
@@
 Limitations.
-Our method suffers from the following limitations: (a) We found our method to be incapable of rotating objects, and instead stretch the object to fit the new blob shape without changing the orientation.
-(b) We found our method to struggle with resizing object, especially in large resizes (e.g., Resize 3 in the second row).
-(c) We found our method to struggle to handle colliding object while dragging, which may result with a hybrid between the objects (e.g., Drag 2 in the third row) or one object being merged (e.g., Drag 3 in the third row).
+Our method suffers from the following limitations: (a) We found our method to be incapable of rotating objects, and instead stretch the objects to fit the new blob shape without changing the orientation.
+(b) We found our method to struggle with resizing objects, especially in large resizes (e.g., Resize 3 in the second row).
+(c) We found our method to struggle to handle colliding objects while dragging, which may result with a hybrid between the objects (e.g., Drag 2 in the third row) or one object being merged (e.g., Drag 3 in the third row).
 Our method suffers from the following limitations that are depicted in Figure 7: (a) We found our diffusion anchoring technique introduced in Section 4.2 to be incapable of rotating objects, and instead, as can be seen in Figure 7(a), stretch the object to fit the new blob shape without changing the orientation, this may be caused due to the fact that rotation involves understanding the 3D structure, which is not reflected by the self-attention nearest-neighbor copying.
@@
 (c) We found our method to struggle to handle colliding object while dragging, which may result with a hybrid between the objects (Figure 7(c) Drag 2) or one object being merged (Figure 7(c) Drag 3).
-In conclusion, we presented DiffUHall, our solution to the seemingly straightforward task of object dragging.
+In conclusion, we presented DiffUHaul, our solution to the seemingly straightforward task of object dragging.
 We demonstrated that the spatial understanding of the localized BlobGEN can be harnessed to this task, using our novel diffusion anchoring technique that manages to merge the location signal from the model with the object appearance signal from the input image.
-We hope that our contributions will serve as a valuable tools not only for object dragging but also for other creative tasks in the future.
+Acknowledgements.
+This work was supported in part by the Israel Science Foundation (grants 2492/20, 3611/21 and 1574/21).
 Input
@@
 Our inference-time masking mechanism (right) constrain the attention between the textual tokens and only their corresponding visual tokens within their blobs.
-As explained in Section 4.1, we offer a method to visualize the gated self-attention layer of GLIGEN (Li et al., 2023) and BlobGEN (Nie et al., 2024).
+As explained in Section 4.1 we offer a method to visualize the gated self-attention layer of GLIGEN (Li et al., 2023) and BlobGEN (Nie et al., 2024).
 In gated self-attention, a projection layer first converts the CLIP (Radford et al., 2021) text embeddings of the text description $S_{i}$ to the text tokens $T=\{t_{1},...t_{n}\}$.
@@
 As can be seen, all these components improve the foreground object consistency.
+For example, see the distorted face of the zebra or the distorted shape of the water pipe in the ablated cases.
 In Figure 13 we provide an additional qualitative comparison of our method against the baselines on the automatically extracted dataset (as explained in Section A.3).
@@
 In addition, in Figure 14 we provide a qualitative visualization of the ablation study we conducted.
-As can be seen, PBE (Yang et al., 2022), DiffusionSG (Epstein et al., 2023) and Anydoor (Chen et al., 2023a) mainly suffer from a bad preservation of the foreground object.
-DragDiffusion (Shi et al., 2023) struggles with dragging the object, while DragonDiffusion (Mou et al., 2023) and DiffEditor (Mou et al., 2024) suffers from object traces.
-Our method, on the other hand, strikes the balance between dragging the object and preserving its identity.
+As can be seen, all these components improve the foreground object consistency.
+Please notice, the contribution of the GSA masking, as reflected by our automatic metric in Table 2 in the main paper, is limited in comparison to the other components that we added.
+However, the GSA masking mitigates the GSA leakage, which affects changes in the identity of the moved objects.
+This effect is visualized in Figure 14: see the distorted face of the zebra or the distorted shape of the water pipe in the second row, which ablates GSA masking.
 Appendix C Societal Impact